What 2-digit number is three times the sum of its digits?
Answer by uma(370)

The two digit number is 27
let the number be 10+y
Sum of the digits = x+y
So 10x+y = 3(x+y)
==> 10x + y - 3x -3y = 0
==> 7x - 2y = 0
On giving x = 2 we find that y = 7
Therefore the required number is 27